OpenStudy (anonymous):

Can this be simplified? Also my as well check to make sure it is correct. Find the orthogonal projection of f onto g. C[-1,1], f(x)=x, g(x)=e^x

OpenStudy (anonymous):

\[<f,g>=\int\limits_{a}^{b}f(x)g(x)dx\]

OpenStudy (anonymous):

\[\frac{2e^{-1}}{\frac{1}{2}(e^2-e^{-2})}\]

OpenStudy (anonymous):

teacher likes to mess around with equations if he doesn't think they are hard enough

OpenStudy (anonymous):

for <f,g> i got 2e^-1

OpenStudy (anonymous):

this is only the constant in front o the vector also

OpenStudy (anonymous):

so my answer above would have an x at the end

OpenStudy (anonymous):

\[\int_{-1}^{1} x*e^xdx\]is it like this?

OpenStudy (anonymous):

If it is like that umm then \[\left| x*e^x - e^x\right|_{-1}^{1}\] \[e - e - (-e^{-1} - e^{-1})\] \[-2*e^{-1}\] \[-\frac{2}{e}\] maybe I am not sure

OpenStudy (anonymous):

i don't belive that is correct

OpenStudy (anonymous):

if you use tabular's you get should get (x-1)e^x

OpenStudy (anonymous):

which is what you have aboe

OpenStudy (anonymous):

\[xe^x-e^2=(x-1)e^x\]

OpenStudy (anonymous):

solving that i get\[(1-1)e^x=0e^1-(-1-1)e^{-1}=-2e^{-1}=2e^{-1}\]

OpenStudy (anonymous):

you are missing a negative ithink somewhere

OpenStudy (anonymous):

oh yeah you are right

OpenStudy (anonymous):

so then you have to do the inner product of <g,g> = just switch f(x)g(x) to g(x)g(x)dx

OpenStudy (anonymous):

and tell me what you get there

OpenStudy (anonymous):

\[\frac{e^{2x}}{2}\] evaluated from -1 to 1

OpenStudy (anonymous):

g(x).g(x) = e^(2x) \[|\frac{e^{2x}}{2}|_{-1}^{1}\] \[\frac{e^2}{2} - \frac{e^{-2}}{2}\]

OpenStudy (anonymous):

is there anyway to simplify that? is there a identity?

OpenStudy (anonymous):

\[e^x+e^{-x}/2=\]

OpenStudy (anonymous):

either sinh cosh? or no?

OpenStudy (anonymous):

simplify that hmm no idea

OpenStudy (anonymous):

\[\sinh(2x) or \cosh(2x)?\]

OpenStudy (anonymous):

it'd be cosh(2x) but i forget how hyperbolic functions work again?

OpenStudy (anonymous):

hmm sorry I haven't even studied hyperbolic functions